OpenStudy (anonymous):

(4x^3 + 3x^2 - 30x - 10) ÷ (x -3)

jhonyy9 (jhonyy9):

3x*3 +3x*2 -30x -10/x-3 = 3x*2+12x +90 3x*3-9x*2 --------- -- 12x*2 -30x 12x*2 -120x ----------- --- 90x-10 90x-270 -------- -- 260

jhonyy9 (jhonyy9):

sorry this begin with 4x*2 and not 3x*2

OpenStudy (anonymous):

thanks!! :)

jhonyy9 (jhonyy9):

3x*3 +3x*2 -30x -10/x-3 = 3x*2+12x +90 3x*3-9x*2 --------- -- 12x*2 -30x 12x*2 -120x ----------- --- 90x-10 90x-270 -------- -- 260

jhonyy9 (jhonyy9):

can you understanding hence too or i rewrite it ?

OpenStudy (anonymous):

no its good

jhonyy9 (jhonyy9):

3x*3 +3x*2 -30x -10/x-3 = 3x*2+12x +90 3x*3-9x*2 --------- -- 12x*2 -30x 12x*2 -120x ----------- --- 90x-10 90x-270 -------- -- 260

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

3x*3 +3x*2 -30x -10/x-3 = 3x*2+12x +90 3x*3-9x*2 --------- -- 12x*2 -30x 12x*2 -120x ----------- --- 90x-10 90x-270 -------- -- 260

jhonyy9 (jhonyy9):

can you give me medal too ?

OpenStudy (anonymous):

yeah, but i dont know how since they changed it

jhonyy9 (jhonyy9):

what do you like changing,sorry ? if the ecuation begin with 4x*3 and not with 3x*3 the result after the equal sign need begining to you not with 3x*2 - this need begining with 4x*2

jhonyy9 (jhonyy9):

because 4x*2 multiplied with x will add to you 4x*3

jhonyy9 (jhonyy9):

in the rest is correct

jhonyy9 (jhonyy9):

hope so much that now it is clearly

OpenStudy (anonymous):

yep thank you

jhonyy9 (jhonyy9):

please click on good answer please